A car first move √2 km due northeast 2 km due north and finally 3km due east .then the total displacement of the car is.

Grade:12th pass

1 Answers

Susmita
5 years ago
Let the car be initially at origin.
Travells north east means its trajectory is y=tan45x
Or,y=x
So x2+x2=(root2)2=2
or,2x2=2
or,x=1. (We would not take x=-1 as it is not moving north west.)
So y=1
So after travelling root2 km north east,the car's coordinate is (1,1).
Now the car moves 2km north.So the x coordinate remains unchanged and y coordinate increases by 2 km.So the car's coordinate after travveling north is (1,3).
Now the car moves 3km east.so y coordinate does not change and x coordinate increases by 3km.so car's final coordinate is (4,3).
So displacement is distance from origin to (4,3) that is
Root(42+32)=root(16+9)=root25=5 km